A car started from Phoenix toward Los Angeles at the rate of 50 mi/h. A second car left from the same place 1 h later and drove along the same route at 65 mi/h. How long did it take the second car to overtake the first?

Answer by mathstutor458(57)

let the time is 't'
A car started from Phoenix toward Los Angeles at the rate of 50 mi/h.
A second car left from the same place 1 h later and drove along the same route at 65 mi/h.
distance travelled by both cars at overtaking point is same.
Therefore, 50(t+1) = 65t
=>50t+50 = 65t
=>15t = 50 =>t = 50/15 =3.33hrs